Subject: APO 3.5m users committee minutes, July 2009
From: Suzanne Hawley
Submitted: Wed, 15 Jul 2009 12:44:02 -0700 (PDT)
Message number: 1139
(previous: 1138,
next: 1140
up: Index)
APO 3.5-m Users Committee Phonecon, 7/13/09
Attending: Suzanne Hawley, Ed Turner (for Michael Strauss), Jon
Fulbright, Mark Klaene, Russet McMillan, John Bally, Bruce Gillespie
Absent: Scott Anderson, Michael Strauss, Al Harper, Jon Holtzman,
Remy Indebetouw
**********************************
User feedback, comments from institutional representatives:
Princeton (Ed Turner) - Ed had not heard of any user comments.
Colorado (John Bally) - No general user comments, but John Stocke
asked if a tip-tilt capability was being considered as part of the
visible-light imager upgrade. Suzanne said that it was not part of
the landscape for the imager upgrade at this point; tip-tilt options
for the telescope and instrument had been considered extensively in
recent years, and the cost-benefit arguments were not compelling. We
are looking at the QUOTA camera from WIYN which uses orthogonal transfer
arrays and provides similar tip-tilt correction. Bruce Woodgate is
studying tip-tilt as a possible option on upgrades to the Goddard
Fabry-Perot, and John said he would contact Bruce to see where
he was.
Johns Hopkins (Jon Fulbright) - Jon had not heard any comments from
JHU users.
Washington (Suzanne Hawley) - Things are basically fine. UW users
are looking into the Agile dome-flats issue.
Chicago (Al Harper) - No report.
New Mexico State (Jon Holtzman) - No report.
Virginia (Remy Indebetouw) - No report.
**********************************
Discussion of telescope/instruments report:
Regarding the more detailed report below, Mark said that weather was a
major factor in observing success recently; we had unusually wet
weather in late June, but the conditions have surprisingly dried out
and observing has resumed. Another activity worth mentioning is the
influx of SDSS-III BOSS people, who are on site making preparations
for the hardware and software upgrades for BOSS during the summer.
For the telescope performance, things have been generally pretty
good. During the June mini-shutdown, which went well, the new
altitude drive controller was installed and tested. Tuning is still
underway but the performance so far is very good, for which Fritz
Stauffer deserves considerable credit. Agile was moved to the
TR2 port with its own dedicated rotator and 3-inch filter wheel.
There is still some cleanup to do, but Agile is operational at its new
permanent station.
For instruments, SPIcam, the echelle, and NIC-FPS are all functioning
well. There have been minor problems with the DIS vacuum system; new
ion pumps have been ordered, and a long-term solution to the DIS
vacuum maladies are under research. The formal acceptance and
handover of TripleSpec is moving forward, albeit slowly. Extensive
work has been done to study stray light issues in Agile (see later
topic, this meeting), and we are modifying its filter mount to accept
2-inch filters. APOLLO observing has been generally fine.
The CIF projects arena is focussed on completion and installation of
the direct drives for altitude and azimuth. The drive assemblies are
in the final phases of machining, and are expected at APO next week,
which is close to the original schedule. The big summer shutdown for
their installation is slated to begin 10 August.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 6/04/09
through 7/7/09
Mark Klaene
0) Overview
The monsoon rains arrived a little early this year, around June 15th--
we have already received four inches of rain for the season.
Naturally, we have experienced more lost observing time because of the
increased cloud cover, precipitation, and high humidity. The SDSS
BOSS group has also arrived en masse to start spectrograph and
software integration; this will continue to tax our on-site housing
resources throughout the summer. The primary and tertiary mirrors
were washed during the short shutdown this week.
1) Telescope
Telescope operations have been generally smooth, with some significant
progress accomplished during the June engineering time. This includes
the installation of the new altitude axis controller, and the move of
Agile to its permanent mount on the TR2 port. The transition to the
new axis controller has gone better than we could have hoped for.
Fritz Stauffer's hard work on this, and the prior integration of the
NA2 and TR2 axes, paid off. This will yield a major improvement in
long-term maintainability of the telescope and eliminates a concern we
have had for years with respect to failure of the old obsolete drive
controllers. In addition, the new controllers are paving the way for
the direct-drive motors with their promise of improved pointing and
tracking performance.
2) Instruments
SPIcam, Echelle and NIC-FPS have been operational with no problems.
The high-frequency pattern noise on SPIcam reported last month seems
to have subsided.
The DIS ion vacuum pumps have given us some trouble during this
period, but we have been able to keep the instrument operational.
For TripleSpec, we are starting to make progress again on the
instrument acceptance documentation. Currently the digital slit mask
is slightly off and we are waiting for available time to generate a
new one.
Agile is operational on TR2. The current status of the dome flats are
under discussion and the instrument baffling is less than ideal--sky
flats are an acceptable alternative. There are a number of small
clean-up items remaining, most of which have to do with operational
issues like filter wheel handling, balance, homing, and zeroing--these
are being addressed as time permits. We also hope to have the 2-inch
filter option back soon.
The APOLLO laser system is operational.
3) CIF projects
The direct-drive project is well on its way to having all of the
telescope parts at the site by late July. Machining schedules are
very tight, but we are still on course for installation during the
summer shutdown, scheduled for Aug 10-Sep 6. The new azimuth drive
controller is being developed and initial testing has begun.
**********************************
Update on recent July engineering work (Klaene):
Mark reported that the APO engineering staff successfully washed M1
and M3, and the new hardware for the altitude and azimuth drive
controllers was installed and tested. All three telescope motions
(alt, az, and rotators) are now running on new controllers, which is a
major accomplishment and removes our dependency and risk related to
the old controllers that were obsolete and unmaintainable. Also, the
M3 rotation logic has been modified so that the tertiary mirror moves
more quickly between the primary science ports. Some improvements to
the tertiary mirror tip-tilt adjustments are being explored.
**********************************
Agile dome flats at the TR2 port (Hawley):
At its new TR2 port, Agile dome flats show significant scattered
light. Sky flats appear to be OK, but on-sky tests near the moon
have not yet been carried out. Recent tests have suggested several
ways to improve the dome flats by identifying sources of scattered
light and changing the instrument baffling. These are under
investigation. We are also considering building a library of
good dome flats and sky flats for Agile users. In the meantime,
users should not use dome flats for science, and should instead obtain
sky flats. We also will appreciate user feedback on their flats
and science data.
**********************************
Improvements in FITS headers (Hawley):
There are 2 issues with FITS headers to discuss. First, the headers
have until recently contained incorrect gain and read noise
parameters for the DIS red chip, being the parameters from the old CCD.
These were replaced with the proper parameters for the new red CCD (which
was installed in Dec 2006) in late June and are now correct in the DIS
FITS headers. Second, LST entries in the headers were not correct
for the start of the exposure time. This bug has apparently existed
for at least the past 10 years. If data has LST written in scientific
notation, it should not be trusted. The LST is now written in
sexagesimal notation (HHMMSS) and is accurate to within a second
of the exposure start time. Note that the UT times written in the
headers have always been correct, and if users need LSTs for older
exposures, they can obtain them via manual computations based on the
other data in the headers. Committee members should make sure that
their users know that these changes in the FITS headers have been made.
We appreciate the feedback from several users recently which led to us
uncovering and fixing these issues.
**********************************
Update on imaging camera and echelle upgrade studies (Hawley):
Cynthia Froning and Eric Burgh (both at CU) are heading up the studies
for an optical imager and echelle spectrograph upgrade, respectively.
Cynthia provided the following update on the imager study:
Chris Burrows has been hired as a consultant and he has completed a
preliminary optical design for a camera that covers a 10 arcmin FOV
with no vignetting at 0.35 arcsec/pixel sampling. We are currently
using this baseline design to evaluate the trade space with regard
to field of view, spatial sampling, and broadband and narrowband
performance versus complexity and cost. Our next steps will be
to extend the study to evaluating the feasibility of placing a camera at
the TR1 port rather than NA2 and examining the cost vs. performance
enhancements of expanding to a camera that provides simultaneous multicolor
imaging in multiple channels. We plan to complete this work over the next
two months.
Suzanne mentioned that the QUOTA camera from WIYN is also under
consideration during the imager study.
Eric Burgh provided this update on the echelle study:
Still hunting down current component level efficiencies for the detector,
grating, and optics. Need these for adequate comparison to new components.
Received quote from Newport for an echelle grating with better
efficiency that could replace the current one. Am still looking into
quotes for AR coatings for the optics, as well as for the CCDs.
Don York is sending a large stack of echelle documentation to APO
which may be helpful. Also, John Bally said that he recently discussed
the echelle with Dale Sanford from Yerkes/UChicago who was involved in the
building of the echelle. He will put Dale and Eric in touch.
**********************************
Shutdown status & plans (Klaene):
Mark reminded the committee that the main summer shutdown starts on 10
August. The main event is to replace the altitude and azimuth drive
systems with direct-drive motors. Users should be aware that with the
usual summer weather vagaries, returning the telescope for science
observing is tricky to schedule--we might be early, on schedule, or
late. Suzanne said that she will monitor the shutdown progress
closely, and will alert the institutional schedulers if there is a
need for programs to add if we finish the shutdown early, and will notify
affected users if we are delayed in finishing the engineering work.
**********************************
Date for next meeting (Hawley):
As we traditionally do, we will skip August. The next Users Committee
telecon will be on 14 September at the usual time (8:30am PDT).
**********************************
Miscellaneous topics:
o Suzanne mentioned that there are several postdocs at UW who need to
have their training orientation trip to APO. If other institutions
have a similar situation, they should contact her to see if a pan-
institutional postdoc-training program could be tried out. Bruce said
such a program should be advertised APO-wide, and he wondered if it
might be possible to block some contiguous nights of telescope time
for this purpose. Suzanne said there are various scheduling
possibilities if we know ahead of time (by the previous quarter)
who/how many would be participating.
o Ed Turner asked the committee if anyone had any comments about
endorsing NMSU's request for ARC participation in the SONG proposal
(SONG is the stellar equivalent of the GONG project). There were no
objections from the attendees.
**********************************
ACTION ITEMS:
[open from previous months]:
none
[new actions from this meeting]:
none
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 14
September (no meeting in August), at 8:30 AM Pacific Time. The agenda
and other materials will be sent to the committee members during the
preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1139 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO